# Mirror symmetry and rigid structures of generalized K3 surfaces

### v5

@@ Abstract.
The present article is concerned with mirror symmetry for generalized K3 surfaces, with particular emphasis on complex and Kähler rigid structures. Inspired by the works of Dolgachev, Aspinwall–Morrison and Huybrechts, we introduce a formulation of mirror symmetry for generalized K3 surfaces by Mukai lattice polarizations, fixing the problems in the conventional formulation of mirror symmetry for K3 surfaces. In particular, we settle the long-standing problem of mirror symmetry for singular K3 surfaces. Along the way, we investigate complex and Kähler rigid structures of generalized K3 surfaces.
@@
mirror symmetry, K3 surfaces, generalized Calabi–Yau manifolds, complex rigidity, Kähler rigidity, B-fields
@@ 1. Introduction
Mirror symmetry for a K3 surface $S$ is a very subtle problem as the complex and Kähler structures are somewhat mixed as they both live in $H^{2}(S,\mathbb{C})$. In the foundational article [3], Dolgachev formulated mirror symmetry for lattice polarized K3 surfaces as a certain duality between the algebraic and transcendental cycles. Although his formulation works nicely in many cases, it cannot be a definitive one as there is an assumption which does not hold in general. An obvious exception is a singular K3 surface, which is a K3 surface with the maximal Picard number $20$. It is also known as a rigid K3 surface in the sense that it admits no deformation of complex structure keeping the maximal Picard number. A satisfactory formulation of mirror symmetry for K3 surfaces, which in particular should solve the puzzle of mirror symmetry for singular K3 surfaces, has been long-anticipated.
The celebrated work [2] of Aspinwall–Morrison is the key articles explaining mirror symmetry for K3 surfaces from a physics point of view. They discussed SCFTs on a K3 surface $S$ and showed that the moduli space $\mathfrak{M}_{(2,2)}$ of $N=(2,2)$ SCFTs fibers over the moduli space $\mathfrak{M}_{(4,4)}$ of $N=(4,4)$ SCFTs. Mathematically such moduli spaces are related with space-like 4-spaces in $H^{*}(S,\mathbb{R})\cong\mathbb{R}^{4,20}$ equipped with the Mukai pairing, and $\mathfrak{M}_{(2,2)}$ and $\mathfrak{M}_{(4,4)}$ are identified with $\mathrm{Gr}^{po}_{2,2}(H^{*}(M,\mathbb{R}))\cong\mathrm{O}(4,20)/(\mathrm{SO}(2)\times\mathrm{SO}(2)\times\mathrm{O}(20))$ and $\mathrm{Gr}^{po}_{4}(H^{*}(M,\mathbb{R}))\cong\mathrm{O}(4,20)/(\mathrm{SO}(4)\times\mathrm{O}(20))$ respectively. Then mirror symmetry is realized as an involution of these moduli spaces. On the other hand, the moduli space $\mathfrak{M}_{\mathrm{HK}}$ of $B$-field shifts of hyperKähler metrics can be identified with an open dense subset of $\mathfrak{M}_{(4,4)}$ by the period map $\mathfrak{per}_{\mathrm{HK}}$. Inspired by this fact, in the beautiful article [6], Huybrechts showed that the counterpart of $\mathfrak{M}_{(2,2)}$ is given by the moduli space $\mathfrak{M}_{\mathrm{gK3}}$ of generalized K3 surfaces, which are the K3 version of the generalized Calabi–Yau structures introduced by Hitchin [5]. The moduli space $\mathfrak{M}_{\mathrm{K3}}\times H^{2}(M,\mathbb{R})$ of $B$-field shifts of K3 surfaces endowed with a Ricci-flat metric are naturally contained in $\mathfrak{M}_{\mathrm{gK3}}$ of real codimension $2$. To summarize, we have the following diagram.
@@
Then, as Huybrechts pointed out, it is natural to expect that points in $\mathfrak{M}_{\mathrm{K3}}$ might be mirror symmetric to points that are no longer in $\mathfrak{M}_{\mathrm{K3}}$. We will show that this is precisely the source of the problems in the conventional formulation of mirror symmetry for K3 surfaces.
There is a good chemistry between Hitchin’s generalized Calabi-Yau structures and mirror symmetry as they both embrace the $A$-structures (symplectic) and $B$-structures (complex) on an equal footing. Inspired by the works of Dolgachev, Aspinwall–Morrison and Huybrechts, we introduce a formulation of mirror symmetry for generalized K3 surfaces (Section 4). The key features of our formulation is twofold, and both are guided by Hitchin’s theory. The first is to extend the scope of lattice polarizations to the Mukai lattice. Mixture of degrees of cycles is indispensable for generalized K3 surfaces. The second is to treat the $A$- and $B$-structures on a completely equal footing. The Néron–Severi lattice and transcendental lattice are defined in the same fashion, and moreover the lattice polarizations are imposed on both.
The main advantage of our formulation is that the elements in $H^{0}(M,\mathbb{Z})$ and $H^{4}(M,\mathbb{Z})$ are no longer special cycles, and we succeed in removing the artificial assumption in Dolgachev’s formulation of mirror symmetry. In fact, the assumption is a reflection of the hyperbolic lattice $U\cong H^{0}(M,\mathbb{Z})\oplus H^{4}(M,\mathbb{Z})$ under mirror symmetry. Our formulation naturally solves the two fundamental problems in Dolgachev’s formulation (Section 4.2).
Along the way, we investigate complex and Kähler rigid structures of generalized K3 surfaces (Section 3). The notion of a complex rigid structure is enhanced to the level of generalized K3 surfaces. On the other hand, the notion of a rigid Kähler structure has been anticipated for a long time from the viewpoint of mirror symmetry. We define such a notion by using a generalized Calabi-Yau structure, which captures a refined integral structure of the Kähler moduli space.
@@
Section 2 provides a brief summary of generalized K3 structures based on Huybrechts’ work [6]. Section 3 investigates the complex and Kähler rigid structures on generalized K3 surfaces. Section 4 is devoted to mirror symmetry for generalized K3 surfaces in comparison with the classical formulation of that for K3 surfaces. The main problems are described in Section 4.2 and they are solved in Sections 4.5 and 4.6.
Acknowledgement
The author is very grateful to Yu-Wei Fan and Kenji Hashimoto for valuable discussions. He would also like to thank the referees for helpful comments and suggestions, especially pointing out [9]. Some parts of this work was done while the author was affiliated by Kyoto University with Leading Initiative for Excellent Young Researchers Grant. This work is partially supported by the JSPS Grant-in-Aid Wakate(B) 17K17817 and Kiban(C) 22K03296.
@@ 2. Generalized K3 surfaces
Hitchin’s invention of generalized Calabi–Yau structures is the key to unify the symplectic and complex structures [5]. Such structures have been extensively studied in 2-dimensions by Huybrechts [6]. In this section, for the sake of completeness, we provide a brief review of Huybrechts’ work [6].
@@ 2.1. Generalized Calabi–Yau structures
$\langle\varphi,\varphi^{\prime}\rangle=\varphi_{2}\wedge\varphi^{\prime}_{2}-\varphi_{0}\wedge\varphi^{\prime}_{4}-\varphi_{4}\wedge\varphi^{\prime}_{0}\in A_{\mathbb{C}}^{4}(M),$
@@
The special appeal of generalized Calabi–Yau manifolds resides in the fact that they embrace complex and symplectic structures on an equal footing. There are two fundamental Calabi–Yau structures:
@@
    $\varphi=e^{\sqrt{-1}\omega}=1+\sqrt{-1}\omega-\frac{1}{2}\omega^{2}.$
@@
$e^{B}\varphi=(1+B+\frac{1}{2}B\wedge B)\wedge\varphi.$
This action is orthogonal with respect to the Mukai pairing, namely $\langle e^{B}\varphi,e^{B}\varphi^{\prime}\rangle=\langle\varphi,\varphi^{\prime}\rangle$. A real closed 2-form is called a $B$-field. For a $B$-field $B$ and a generalized Calabi–Yau structure $\varphi$, the $B$-field transform $e^{B}\varphi$ is a generalized Calabi–Yau structure.
We will later show that a $B$-field is indispensable when we view complex and symplectic structures as special instance of a more general notion. The following shows that a generalized Calabi–Yau structure is a $B$-field transform of either one of the fundamental Calabi–Yau structures.
@@
    with is a holomorphic 2-form $\sigma$ (for a complex structure) and a $B$-field $B$. $B^{0,2}$ is the $(0,2)$-part of $B$ with respect to $\sigma$.
@@
The most fascinating aspects of generalized Calabi–Yau structures is the occurrence of the classical Calabi–Yau structure $\sigma$ (type $B$) as well as of symplectic generalized Calabi–Yau structure $e^{\sqrt{-1}\omega}$ (type $A$) in the same moduli space. This allows us to pass from the symplectic to the complex world in a continuous fashion.
@@
$\varphi_{t}=te^{\frac{1}{t}(\mathrm{Re}(\sigma)+\sqrt{-1}\mathrm{Im}(\sigma))}$
@@
$P_{\varphi}$ $=\mathbb{R}\mathrm{Re}\varphi\oplus\mathbb{R}\mathrm{Im}\varphi\subset A^{*}(M),$
$P_{[\varphi]}$ $=\mathbb{R}[\mathrm{Re}\varphi]\oplus\mathbb{R}[\mathrm{Im}\varphi]\subset H^{*}(M,\mathbb{R})$
i.e. the vector spaces spanned by the real and imaginary parts of $\varphi$ and $[\varphi]$ respectively.
$P_{\varphi}$ is positive with respect to the pairing $\langle$-,$\rangle$ - , - ⟩ at every point. $P_{[\varphi]}\subset H^{*}(M,\mathbb{R})\cong\mathbb{R}^{4,20}$ is a positive plane with respect to the Makai pairing.
We say that generalized Calabi–Yau structures $\varphi$ and $\varphi^{\prime}$ are orthogonal if $P_{\varphi}$ and $P_{\varphi^{\prime}}$ are pointwise orthogonal. This orthogonality is in general a stronger condition than just $\langle\varphi,\varphi^{\prime}\rangle=0$.
@@
Let $\varphi=\sigma$ a holomorphic $2$-form for a complex structure on $M_{\sigma}$. If $\varphi^{\prime}$ is a Kähler structure for $\varphi$, then it is of the form $\varphi^{\prime}=\varphi^{\prime}_{0}e^{B+\sqrt{-1}\omega}$ (type $A$) for some $B$-field $B$ and symplectic form $\omega$. The orthogonality is equivalent to $\sigma\wedge B=\sigma\wedge\omega=0$. Therefore $B$ is a closed real $(1,1)$-form and either ±$\pm\omega$ is a Kähler form with respect to $\sigma$.
@@
A generalized Calabi–Yau structure $\varphi$ is hyperKähler if there exists a Kähler structure $\varphi^{\prime}$ such that $\langle\varphi,\overline{\varphi}\rangle=\langle\varphi^{\prime},\overline{\varphi^{\prime}}\rangle$. Such $\varphi^{\prime}$ is called a hyperKähler structure for $\varphi$.
@@
Any hyperKähler structure is a $B$-field transform of one of the above cases. In summary, there are only 3 possible combinations: (type $A$, type $B$), (type $B$, type $A$), (type $A$, type $A$).
@@ 2.3. Generalized K3 surfaces
A $B$-field transform $e^{B}(\varphi,\varphi^{\prime})=(e^{B}\varphi,e^{B}\varphi^{\prime})$ of a generalized K3 surface $(\varphi,\varphi^{\prime})$ is a generalized K3 surface. Generalized K3 surfaces $(\varphi,\varphi^{\prime})$ and $(\psi,\psi^{\prime})$ are called isomorphic if there exist a diffeomorphism $f\in\mathrm{Diff}_{*}(M)$ and an exact $B$-field $B\in A^{2}(M)$ such that $(\varphi,\varphi^{\prime})=e^{B}f^{*}(\psi,\psi^{\prime})$. As indicated in the introduction, the moduli space $\mathfrak{M}_{\mathrm{gK3}}$ of generalized K3 surfaces has a nice description in terms of cohomology groups. Namely, the period map
@@
is an immersion with dense image. Here $\mathrm{Gr}^{po}_{2,2}(H^{*}(M,\mathbb{R}))$ is the Grassmannian parametrizing the orthogonal pairs of positive oriented 2-planes in $H^{*}(M,\mathbb{R})\cong\mathbb{R}^{4,20}$. The details are lengthy to present here and we refer the reader to [6].
2.4. Moduli spaces and Torelli type theorem
We lastly compare the moduli space $\mathfrak{N}_{\mathrm{gCY}}=\{\mathbb{C}\varphi\}/\cong$ of the generalized Calabi–Yau structures of hyperKähler type, and the moduli space $\mathfrak{N}_{\mathrm{K3}}=\{\mathbb{C}\sigma\}/\mathrm{Diff}_{*}(M)$ of the classical marked K3 surfaces.
Theorem 2.10 ([6]).
The classical period map $\mathfrak{per}_{\mathrm{K3}}$ naturally extends to the period map $\mathfrak{per}_{\mathrm{gCY}}$ for the generalized Calabi–Yau structures of hyperKähler type.

### v6

@@ Abstract.
The present article is concerned with mirror symmetry for generalized K3 surfaces, with particular emphasis on complex and Kähler rigid structures. Inspired by the works of Dolgachev, Aspinwall–Morrison and Huybrechts, we introduce a formulation of mirror symmetry for generalized K3 surfaces by using Mukai lattice polarizations. This approach solves issues in the conventional formulations of mirror symmetry for K3 surfaces. In particular, we provide a solution to the problem of mirror symmetry for singular K3 surfaces. Along the way, we investigate complex and Kähler rigid structures of generalized K3 surfaces.
@@
mirror symmetry, K3 surfaces, generalized Calabi–Yau manifolds, complex rigidity, Kähler rigidity, non-commutative geometry
@@ 1. Introduction
Mirror symmetry for a K3 surface $S$ is a very subtle problem, as the complex and Kähler structures are somewhat mixed in $H^{2}(S,\mathbb{C})$. In the foundational article [7], Dolgachev formulated mirror symmetry for lattice polarized K3 surfaces as a duality between the algebraic and transcendental cycles. Although his formulation works well in many cases, it cannot be definitive due to an assumption that does not hold in general. A notable exception is a singular K3 surface, which has the maximal Picard number $20$. This kind of K3 surfaces, also known as a rigid K3 surface, admits no deformation of complex structure while retaining the maximal Picard number. A satisfactory formulation of mirror symmetry for K3 surfaces, particularly one that solves the puzzle of mirror symmetry for singular K3 surfaces, has been anticipated.
The celebrated work [3] of Aspinwall–Morrison is a key article explaining mirror symmetry for K3 surfaces from a physics perspective. They discussed SCFTs on a K3 surface $S$ and demonstrated that the moduli space $\mathfrak{M}_{(2,2)}$ of $N=(2,2)$ SCFTs fibers over the moduli space $\mathfrak{M}_{(4,4)}$ of $N=(4,4)$ SCFTs. Mathematically such moduli spaces are related with space-like 4-spaces in $H^{*}(S,\mathbb{R})\cong\mathbb{R}^{4,20}$ equipped with the Mukai pairing, and $\mathfrak{M}_{(2,2)}$ and $\mathfrak{M}_{(4,4)}$ are identified with certain Grassmannians $\mathrm{Gr}^{po}_{2,2}(H^{*}(M,\mathbb{R}))\cong\mathrm{O}(4,20)/(\mathrm{SO}(2)\times\mathrm{SO}(2)\times\mathrm{O}(20))$ and $\mathrm{Gr}^{po}_{4}(H^{*}(M,\mathbb{R}))\cong\mathrm{O}(4,20)/(\mathrm{SO}(4)\times\mathrm{O}(20))$ respectively. Then mirror symmetry is realized as an involution of these moduli spaces. On the other hand, the moduli space $\mathfrak{M}_{\mathrm{HK}}$ of $B$-field shifts of hyperKähler metrics can be identified with an open dense subset of $\mathfrak{M}_{(4,4)}$ by the period map $\mathfrak{per}_{\mathrm{HK}}$. Inspired by this fact, Huybrechts showed in [13] that the counterpart of $\mathfrak{M}_{(2,2)}$ is given by the moduli space $\mathfrak{M}_{\mathrm{gK3}}$ of generalized K3 surfaces, which are the K3 version of the generalized Calabi–Yau structures introduced by Hitchin [12]. The moduli space $\mathfrak{M}_{\mathrm{K3}}\times H^{2}(M,\mathbb{R})$ of $B$-field shifts of K3 surfaces endowed with a Ricci-flat metric is naturally contained in $\mathfrak{M}_{\mathrm{gK3}}$ of real codimension $2$. To summarize, there is the following diagram.
@@
It is therefore natural to expect that points in $\mathfrak{M}_{\mathrm{K3}}$ might be mirror symmetric to points that are no longer in $\mathfrak{M}_{\mathrm{K3}}$. We will demonstrate that this is precisely the source of the problems in the conventional formulation of mirror symmetry for K3 surfaces (cf. Remark 4.6).
There is a good chemistry between Hitchin’s generalized Calabi–Yau structures and mirror symmetry as they both embrace the symplectic and complex on an equal footing. Inspired by works of Dolgachev, Aspinwall–Morrison and Huybrechts, we introduce a formulation of mirror symmetry for generalized K3 surfaces (Section 4). The key features of our formulation are twofold, and both are guided by Hitchin’s theory. The first is to extend the scope of lattice polarizations to the Mukai lattice. Mixture of degrees of cycles is indispensable for generalized K3 surfaces. The second is to treat the two structures in mirror symmetry on completely equal footing. The Néron–Severi lattice and transcendental lattice are defined in the same fashion, and moreover the lattice polarizations are imposed on both.
The primary advantage of our formulation is that the elements in $H^{0}(M,\mathbb{Z})$ and $H^{4}(M,\mathbb{Z})$ are no longer treated as special. This allows us to eliminate the artificial assumption present in Dolgachev’s formulation of mirror symmetry. Indeed, the assumption is a reflection of the existence of the hyperbolic lattice $U\cong H^{0}(M,\mathbb{Z})\oplus H^{4}(M,\mathbb{Z})$ under mirror symmetry. Our formulation naturally settles the two fundamental problems in the conventional formulation (Section 4.2).
Along the way, we investigate the complex and Kähler rigid structures of generalized K3 surfaces (Section 3). The notion of a complex rigid structure is enhanced to the level of generalized K3 surfaces. On the other hand, the notion of a Kähler rigid structure naturally appears as a counterpart of complex rigid structure from the viewpoint of mirror symmetry. We define this notion by using a generalized Calabi–Yau structure, which captures an interesting sublattice of $H^{*}(M,\mathbb{Z})$ derived from the Kähler moduli space (Section 3.4).
@@
Section 2 provides a brief summary of generalized K3 structures based on Huybrechts’ work [13]. Section 3 investigates the complex and Kähler rigid structures on generalized K3 surfaces. Section 4 focuses on mirror symmetry for generalized K3 surfaces in comparison with the classical formulation of that for K3 surfaces. The main problems are described in Section 4.2 and they are solved in Section 4.5 and Section 4.6.
Acknowledgment
The author sincerely thanks Emma Brakkee, Yu-Wei Fan, Kenji Hashimoto and Shinobu Hosono for their valuable discussions. Special gratitude is also extended to the referees for their helpful comments and suggestions, particularly regarding the references to [18]. This work is supported by the JSPS Grant-in-Aid Wakate(B) 17K17817, Kiban(C) 22K03296 and Kiban(A) 23H00083.
@@ 2. Generalized K3 surfaces
Hitchin’s invention of generalized Calabi–Yau structures is a key to unify the symplectic and complex structures [12]. Such structures have been extensively studied in 2-dimensions by Huybrechts [13]. In this section, for the sake of completeness, we provide a brief summary of Huybrechts’ work.
@@ 2.1. Generalized Calabi–Yau structures
$\langle\varphi,\varphi^{\prime}\rangle:=\varphi_{2}\wedge\varphi^{\prime}_{2}-\varphi_{0}\wedge\varphi^{\prime}_{4}-\varphi_{4}\wedge\varphi^{\prime}_{0}\in A_{\mathbb{C}}^{4}(M),$
@@
The special appeal of generalized Calabi–Yau structures resides in the fact that they embrace complex and symplectic structures on an equal footing. There are two fundamental Calabi–Yau structures:
@@
    $\varphi=e^{\sqrt{-1}\omega}:=1+\sqrt{-1}\omega-\frac{1}{2}\omega^{2}.$
@@
$e^{B}\varphi:=(1+B+\frac{1}{2}B\wedge B)\wedge\varphi.$
This action is orthogonal with respect to the pairing $\langle$-,$\rangle$ - , - ⟩, namely
$\langle e^{B}\varphi,e^{B}\varphi^{\prime}\rangle=\langle\varphi,\varphi^{\prime}\rangle.$
A real closed 2-form is called a $B$-field. For a $B$-field $B$ and a generalized Calabi–Yau structure $\varphi$, the $B$-field shift $e^{B}\varphi$ is a generalized Calabi–Yau structure.
We will later show that a $B$-field is indispensable when we view complex and symplectic structures as special instance of a more general notion (Remark 2.3). The following shows that a generalized Calabi–Yau structure is a $B$-field shift of either one of the fundamental Calabi–Yau structures.
@@
    with a holomorphic 2-form $\sigma$ (for a complex structure) and a $B$-field $B$. $B^{0,2}$ is the $(0,2)$-part of $B$ with respect to $\sigma$.
@@
The most fascinating aspect of generalized Calabi–Yau structures is the occurrence of the classical Calabi–Yau structure $\sigma$ (type $B$) as well as of symplectic generalized Calabi–Yau structure $e^{\sqrt{-1}\omega}$ (type $A$) in the same moduli space. This allows us to pass from the symplectic to the complex world in a continuous fashion.
@@
$\varphi_{t}:=te^{\frac{1}{t}(\mathrm{Re}(\sigma)+\sqrt{-1}\mathrm{Im}(\sigma))}=t+\sigma$
@@
$P_{\varphi}$ $:=\mathbb{R}\mathrm{Re}\varphi\oplus\mathbb{R}\mathrm{Im}\varphi\subset A^{*}(M),$
$P_{[\varphi]}$ $:=\mathbb{R}[\mathrm{Re}\varphi]\oplus\mathbb{R}[\mathrm{Im}\varphi]\subset H^{*}(M,\mathbb{R})$
i.e. the $\mathbb{R}$-vector spaces spanned by the real and imaginary parts of $\varphi$ and $[\varphi]$ respectively.
$P_{\varphi}$ is positive with respect to the pairing $\langle$-,$\rangle$ - , - ⟩ at every point. $P_{[\varphi]}\subset H^{*}(M,\mathbb{R})\cong\mathbb{R}^{4,20}$ is a positive 2-plane with respect to the Makai pairing.
We say that generalized Calabi–Yau structures $\varphi$ and $\varphi^{\prime}$ are orthogonal if $P_{\varphi}$ and $P_{\varphi^{\prime}}$ are pointwise orthogonal. This orthogonality is a stronger condition than just $\langle\varphi,\varphi^{\prime}\rangle=0$.
@@
Example 2.6.
Let $\varphi=\sigma$ be a holomorphic $2$-form for a complex structure on $M_{\sigma}$. If $\varphi^{\prime}$ is a Kähler structure for $\varphi$, then it is of the form $\varphi^{\prime}=\varphi^{\prime}_{0}e^{B+\sqrt{-1}\omega}$ (type $A$) for some scalar $\varphi_{0}^{\prime}$, $B$-field $B$ and symplectic form $\omega$. The orthogonality is equivalent to $\sigma\wedge B=\sigma\wedge\omega=0$. Therefore $B$ is a closed real $(1,1)$-form and either ±$\pm\omega$ is a Kähler form with respect to $\sigma$.
@@
A generalized Calabi–Yau structure $\varphi$ is hyperKähler if there exists a Kähler structure $\varphi^{\prime}$ for $\varphi$ such that
$\langle\varphi,\overline{\varphi}\rangle=\langle\varphi^{\prime},\overline{\varphi^{\prime}}\rangle.$
Such $\varphi^{\prime}$ is called a hyperKähler structure for $\varphi$.
@@
Any hyperKähler structure is a $B$-field shift of one of the above cases. In summary, there are only 3 possible combinations:
(type $A$, type $B$), (type $B$, type $A$), (type $A$, type $A$).
@@ 2.3. Generalized K3 surfaces
A $B$-field shift $e^{B}(\varphi,\varphi^{\prime})=(e^{B}\varphi,e^{B}\varphi^{\prime})$ of a generalized K3 surface $(\varphi,\varphi^{\prime})$ is a generalized K3 surface. Generalized K3 surfaces $(\varphi,\varphi^{\prime})$ and $(\psi,\psi^{\prime})$ are called isomorphic if there exist a diffeomorphism $f\in\mathrm{Diff}_{*}(M)$ and an exact $B$-field $B\in A^{2}(M)$ such that $(\varphi,\varphi^{\prime})=e^{B}f^{*}(\psi,\psi^{\prime})$. As indicated in the introduction, the moduli space $\mathfrak{M}_{\mathrm{gK3}}$ of generalized K3 surfaces has a nice description in terms of the cohomology groups.
Theorem 2.11.
The period map
@@
is an immersion with dense image. Here $\mathrm{Gr}^{po}_{2,2}(H^{*}(M,\mathbb{R}))$ is the Grassmannian parametrizing the orthogonal pairs of positive oriented 2-planes in $H^{*}(M,\mathbb{R})\cong\mathbb{R}^{4,20}$.
A point in the complement of the image of $\mathfrak{per}_{\mathrm{gK3}}$ is orthogonal to a $(-2)$-class and can be interpreted as a degenerate structure.
2.4. Moduli spaces and period domains
We lastly compare the moduli space $\mathfrak{N}_{\mathrm{gCY}}:=\{\mathbb{C}\varphi\}/\cong$ of the generalized Calabi–Yau structures of hyperKähler type, and the moduli space $\mathfrak{N}_{\mathrm{K3}}:=\{\mathbb{C}\sigma\}/\mathrm{Diff}_{*}(M)$ of the classical marked K3 surfaces. (Figure 1)
Theorem 2.12.
The classical period map $\mathfrak{per}_{\mathrm{K3}}$ naturally extends to the period map $\mathfrak{per}_{\mathrm{gCY}}$ for the generalized Calabi–Yau structures of hyperKähler type: